# Accelerating Containerized Service Delivery at the Network Edge

### v1

@@ V Related Works
More recently, Starlight [11] has introduced a novel approach by redesigning the container image architecture. It transmits only the necessary components, significantly reducing transfer times. However, Starlight’s reliance on specific container engines limits its applicability, and it continues to use a traditional client-server registry architecture, which restricts throughput under high bandwidth loads. Future work could integrate Starlight with PeerSync to achieve (a) distributed transmission of large data chunks and (b) selective transmission of necessary data, further improving the containerization ecosystem.
VI Concluding Remarks
In this paper, we introduced PeerSync, an innovative fully decentralized image distribution system tailored for containerized model inference workloads at the network edge. PeerSync leverages advanced P2P downloading techniques, outperforming traditional methods by dynamically adapting to fluctuating network conditions and content popularity. Its embedded autonomous tracker enhances resilience by eliminating single points of failure, making it particularly effective in edge environments characterized by unstable networks and resource-constrained devices. Additionally, the cache cleaner efficiently optimizes storage usage while maintaining performance, addressing a key limitation of edge devices.
Future work offers several opportunities to extend PeerSync’s capabilities. a) Dynamic block sizing: While the current static block size approach performs well, incorporating adaptive block sizes with multiple Merkle tree versions could further enhance efficiency with minimal overhead. b) Support for heterogeneous LANs: Adapting to diverse network profiles, such as WLANs, using a combination of user-defined tags and auto-detection techniques could improve performance in non-uniform environments. c) Plugin-based algorithm design: Providing users with greater flexibility through plugin-based customization of algorithms could further expand PeerSync’s adaptability.
Integrating PeerSync with Starlight [11] represents another exciting avenue for exploration. The combination of PeerSync’s decentralized, adaptive framework and Starlight’s innovative protocols, filesystem, and image storage formats could pave the way for next-generation containerization ecosystems, optimized for the evolving model inference demands at the network edge.

### v2

@@ VI Related Works
More recently, Starlight [10] has introduced a novel approach by redesigning the container image architecture. It transmits only the necessary components, significantly reducing transfer times. However, Starlight’s reliance on specific container engines limits its applicability, and it continues to use a traditional client-server registry architecture, which restricts throughput under high bandwidth loads. Future work could integrate Starlight with PeerSync to achieve distributed transmission of large data chunks and selective transmission of necessary data, further improving the containerization ecosystem.
VII Concluding Remarks
In this paper, we introduced PeerSync, a fully decentralized image distribution system tailored for containerized model inference at the network edge. PeerSync leverages P2P downloading to dynamically adapt to changing network conditions and content popularity, significantly outperforming traditional approaches. Its autonomous tracker eliminates single points of failure, enhancing resilience in unstable and resource-constrained edge environments. The integrated cache cleaner also ensures efficient storage use without compromising performance.
Building on the modular approach, future work could also incorporate intelligent caching as on-demand modules. This would allow for dynamic adaptation to various edge site designs, for example, by implementing proactive caching in predictable environments where rich metrics are available. Another promising direction is integrating PeerSync with Starlight [10]. Combining PeerSync’s P2P distribution with Starlight’s partial transmission and novel image format could enable a next-generation container ecosystem optimized for edge model inference.